OpenStudy (anonymous):

Solve –2(2x – 5) = 2

OpenStudy (anonymous):

-4x+10=2 -4x=-8 x=2

OpenStudy (anonymous):

@sydneygill14 is right

OpenStudy (anonymous):

@sydneygill14 How did you get -4x+10=2 ? I don't see it.

OpenStudy (anonymous):

you have to distribute the -2 throughout the (2x-5) (-2*2x)-(-2*5)=2 -4x+10-2